OpenStudy (mayankdevnani):

POWER OF ENGINEERING ! NOTE :- Don't take this post so seriously .Everything is wrong in it. It's for fun. Enjoy

OpenStudy (mayankdevnani):

Prove \(\large \bf \frac{0}{0}=2\)

OpenStudy (mayankdevnani):

Solution :- \[\large \bf \frac{0}{0}=\frac{100-100}{100-100}\] \[\large \bf =\frac{10.10-10.10}{10.10-10.10}\] \[\large \bf =\frac{10^2-10^2}{10(10-10)}\] \[\large \bf =\frac{(10+10)(10-10)}{10(10-10)}\] \[\large \bf =\frac{10+10}{10}\] \[\large \bf =\frac{20}{10}\] \[\large \bf =2\]

OpenStudy (mayankdevnani):

Proved.

OpenStudy (moldybubblegum12):

wow

OpenStudy (mayankdevnani):

Can anyone tell me where's the mistake in that proof ?

OpenStudy (acespeedfighter):

hmmm let me think

OpenStudy (skullpatrol):

division by 0

OpenStudy (skullpatrol):

you can't just " cancel" 10-10 from the numerator and the denominator

OpenStudy (mayankdevnani):

Cool :D

OpenStudy (skullpatrol):



OpenStudy (skullpatrol):

10 - 10 = 0

OpenStudy (skullpatrol):